OpenStudy (anonymous):

Use the given endpoint R and midpoint M of RS to find the coordinates of the other endpoint S. R(8,0), M(4,-5) --Explain please!

OpenStudy (mathstudent55):

If you are given the coordinates of two points, do you know how to find the midpoint?

OpenStudy (mathstudent55):

The midpoint of the segment with endpoints \( (x_1, y_1) \) and \( (x_2, y_2) \) is \( \left( \dfrac{x_1 +x_2}{2}, \dfrac{y_1 + y_2}{2} \right) \)